OpenStudy (anonymous):

Please please help <3 Nakim simplified 3√2x+x√8x-5√18x got -10x√2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (ash2326):

first let's simplify the expression \[3 \sqrt(2x)+x \sqrt(8x)-5 \sqrt(18x)\] 8=2^2*2 18=3^2*2 so the expression is now \[3 \sqrt(2x)+x \sqrt(4*2x)-5 \sqrt(9*2x)\] \[=\sqrt(2x)+2x \sqrt(2x)-15\sqrt(2x)\] Part A: nakim would have missed the x in the middle term, and had combined all to get -10 sqrt(2x) but the solution is 2 sqrt(2x)(x-6)

OpenStudy (ash2326):

2 (sqrt(2x))(x-6)